# Christian Ghiglino    Alastair Langtry

### v4

@@ B.1 Proof
The proof here follows 1 fairly closely. To make the derivations less cumbersome, define $z\equiv\frac{1}{w}-\alpha$. Now consider the marginal utility of consumption:
$\frac{du_{j}}{dx_{j}}$ $=\gamma Y(\theta_{j})^{-1}+\frac{1}{2}\beta(x_{j}-Y_{j})^{-0.5}-z$
@@
Notice that $\frac{1}{2}\beta(x_{j}-Y_{j})^{-0.5}\geq 0$ whenever it is well-defined. So we must have $\gamma Y(\theta_{j})^{-1}-z\leq 0$ for a best response to exist. This implies that we must have $Y(\theta_{j})\geq\gamma z^{-1}$. Given this, we can now consider $j$’s best response.
$(x_{j}-Y_{j})^{-0.5}$ $=\frac{2}{\beta}z-\gamma Y(\theta_{j})^{-1}$
@@
$\implies x_{j}-\sum_{k}\widehat{G}_{jk}x_{k}$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}$
Since we must have $Y(\theta_{j})\geq\gamma z^{-1}$ (argued above), the Right-Hand Side of the equation above is weakly positive. This implies that $x_{j}^{*}>Y_{j}$, ensures that preferences are well-defined in equilibrium. We now switch to vector notation (in an abuse of notation we use the same notation for a vector of constants as for the equivalent scalar).
@@
$\mathbf{x}-\widehat{G}\mathbf{x}$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2},$
@@
$\mathbf{x}$ $=(I-\widehat{G})^{-1}\left[\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\right].$
@@
We have assumed $|\lambda_{1}|<1$, so $(I-\widehat{G})$ is invertible (Ballester et al. [2006, Theorem 1]). Revering to scalar notation:
@@
$x_{j}=\sum_{k}(I-\widehat{G})^{-1}_{jk}\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}.$
@@
A comment: by construction, $\widehat{G}$ is a block matrix with off-diagonal blocks being all zero. This is because an agent $j$ can only have a link to agent $k$ if $j$ and $k$ share the same identity (i.e. if $\theta_{j}=\theta_{k}$). An important consequence of this is that $\theta_{k}=\theta_{j}$ whenever $[(I-\widehat{G})^{-1}]_{jk}>0$. This means we can treat the term $Y(\theta_{k})^{-1}$ as a constant an pull it out of the sum. And we can use the (standard) definition of Bonacich centrality. So
@@
$x_{j}^{*}=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\mathbb{C}_{j}^{b}$
This would give equilibrium actions for a fixed value of the prototype. But $Y(\theta_{j})$ is endogenous. Recall that by definition $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}$. Hence, by substituting our equation for consumption into this definition:
$Y(\theta_{j})$ $=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\mathbb{C}_{j}^{b}$
$\implies Y(\theta_{j})$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\bar{\mathbb{C}}^{b}_{\theta_{j}},$
$\implies\frac{4}{\beta^{2}}\frac{Y(\theta_{j})}{\bar{\mathbb{C}}^{b}_{\theta_{j}}}$ $=\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}$
using the fact that ¯$\bar{\mathbb{C}}^{b}_{\theta_{j}}$ is the average of $\mathbb{C}_{k}^{b}$ for agents $k\in I(\theta_{j})$. The LHS of the equation above is clearly monotonically increasing in $Y(\theta_{j})$. And in the region $Y(\theta_{j})\geq\gamma z^{-1}$, the RHS is monotonically decreasing in $Y(\theta_{j})$. [^28] This is less obvious, but follows from the fact that (1) $z-\gamma Y(\theta_{j})^{-1}\geq 0$, (2) this term is increasing in $Y(\theta_{j})$, and (3) this term is raised to a negative exponent. So if a solution that satisfies $Y(\theta_{j})\geq\gamma z^{-1}$ exists, it will be unique. Now rearrange terms:
$\frac{\beta^{2}}{4}\cdot\frac{\bar{\mathbb{C}}^{b}_{\theta_{j}}}{Y(\theta_{j})}$ $=z^{2}-2\gamma z\frac{1}{Y(\theta_{j})}+\frac{\gamma^{2}}{Y(\theta_{j})^{2}}$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}Y(\theta_{j})$ $=z^{2}Y(\theta_{j})^{2}-2\gamma zY(\theta_{j})+\gamma^{2}$
$\implies 0$ $=z^{2}Y(\theta_{j})^{2}-\left(2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\right)Y(\theta_{j})+\gamma^{2}$
Then solve for $Y(\theta_{j})$ using the standard quadratic formula.
$Y(\theta_{j})^{=}\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\pm\sqrt{\left(2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\right)^{2}-4z^{2}\gamma^{2}}\right]$
$Y(\theta_{j})^{=}\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\pm\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$
Aside: now we show that the smaller of the two roots is not permitted, because it is less that $\gamma z^{-1}$ (which we have shown above it not possible). Suppose not:
$\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}-\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$ $>\gamma z^{-1}$
$2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}-\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}$ $>2\gamma z$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}$ $>\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}$ $>\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\psi,$
for some $\psi>0$. Contradiction. So we ignore the smaller root. Therefore:
$Y(\theta_{j})=\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$
Finally, we substitute this express for $Y(\theta_{j})$ back into our expression for $x_{j}^{*}$. This yields:
$x_{j}^{*}=\frac{1}{4}\left(\frac{\beta}{z}\right)^{2}\mathbb{C}_{j}^{b}\left(1-\frac{2\gamma z}{2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}}\right)^{-2}.$

### v7

@@ C.1 Proof
The proof here follows 1 fairly closely. Consider the marginal utility of consumption:
$\frac{du_{j}}{dx_{j}}$ $=\alpha-\gamma Y_{\theta_{j}}+\frac{1}{2}\beta(x_{j}-R_{j})^{-0.5}-\frac{1}{w}$
@@
Notice that $\frac{1}{2}\beta(x_{j}-R_{j})^{-0.5}\geq 0$ whenever it is well-defined. So we must have $\gamma Y_{\theta_{j}}-\alpha+1/w>0$ for a best response to exist. This implies that we must have $Y_{\theta_{j}}\geq(\alpha-1/w)/\gamma$. Given this, we can now consider $j$’s best response.
$(x_{j}-R_{j})^{-0.5}$ $=\frac{2}{\beta}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)$
@@
$\implies x_{j}-\sum_{k}\widehat{G}_{jk}x_{k}$ $=\frac{\beta^{2}}{4}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)^{-2}$
We now switch to vector notation (in an abuse of notation we use the same notation for a vector of constants as for the equivalent scalar).
@@
$\mathbf{x}-\widehat{G}\mathbf{x}$ $=\frac{\beta^{2}}{4}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)^{-2},$
@@
$\mathbf{x}$ $=(I-\widehat{G})^{-1}\left[\frac{\beta^{2}}{4}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)^{-2}\right].$
@@
We have assumed $|\lambda_{1}|<1$, so (IG^)(I-$\widehat${G}) is invertible (Ballester et al. [2006, Theorem 1]). Reverting to scalar notation:
@@
$x_{j}=\sum_{k}(I-\widehat{G})^{-1}_{jk}\frac{\beta^{2}}{4}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)^{-2}.$
@@
A comment: by construction, $\widehat{G}$ is a block matrix with off-diagonal blocks being all zero. This is because an agent $j$ can only have a link to agent $k$ if $j$ and $k$ share the same identity (i.e. if $\theta_{j}=\theta_{k}$). An important consequence of this is that $\theta_{k}=\theta_{j}$ whenever $(I-\widehat{G})^{-1}]_{jk}>0$. This means we can treat the term $Y_{\theta_{k}}^{-1}$ as a constant and pull it out of the sum. And we can use the (standard) definition of Bonacich centrality. So
$x_{j}^{*}=\frac{\beta^{2}}{4}\left(\gamma Y_{\theta_{j}}-\alpha+\frac{1}{w}\right)^{-2}C_{j}^{bon}$
Recall that $Y_{\theta_{j}}=\frac{\mathbf{x}_{A}}{\mathbf{x}_{B}}$ if $\theta_{j}=A$ (and similarly if $\theta_{j}=B)$ and so is endogenous, where $\mathbf{x}_{\theta}=\frac{1}{|I(\theta)|}\sum_{k\in I(\theta)}x_{k}$. In an abuse of notation, we will use $Y_{A}$ to denote $Y_{\theta_{j}}$ for $j$ in identity $A$ (and similarly for $Y_{B}$). We will also do the rest of the proof for some $j$ in identity $A$. It is clear that everything works identically for identity $B$. Hence, by substituting our equation for consumption into this definition:
@@
$Y_{A}\quad=\quad\frac{\frac{1}{|I(A)|}\sum_{k\in I(A)}\frac{\beta^{2}}{4}\left(\gamma Y_{A}-\alpha+\frac{1}{w}\right)^{-2}C_{j}^{bon}}{\frac{1}{|I(B)|}\sum_{k\in I(B)}\frac{\beta^{2}}{4}\left(\gamma Y_{B}-\alpha+\frac{1}{w}\right)^{-2}C_{k}^{bon}}\quad=\quad\frac{\left(\gamma Y_{A}-\alpha+\frac{1}{w}\right)^{-2}\mathcal{C}_{A}^{bon}}{\left(\gamma Y_{B}-\alpha+\frac{1}{w}\right)^{-2}\mathcal{C}_{B}^{bon}}.$
where $\mathcal{C}_{\theta}^{bon}$ denotes the average Bonacich centrality of agents in identity group $\theta$. It is also clear from the definition that $Y_{A}=1/Y_{B}$. Substituting this, and rearranging yields:
$\left(\gamma Y_{A}-\alpha+\frac{1}{w}\right)Y_{A}^{0.5}=\left(\gamma\frac{1}{Y_{A}}\alpha+\frac{1}{w}\right)\cdot\left(\frac{\mathcal{C}_{A}^{bon}}{\mathcal{C}_{B}^{bon}}\right)^{0.5},$ (C.2)
which further re-arranges to
$\gamma Y_{A}^{\frac{5}{2}}+\left(\frac{1}{w}-\alpha\right)Y_{A}^{\frac{3}{2}}-\left(\frac{1}{w}-\alpha\right)\left(\frac{\mathcal{C}_{A}^{bon}}{\mathcal{C}_{B}^{bon}}\right)^{0.5}Y-\gamma\left(\frac{\mathcal{C}_{A}^{bon}}{\mathcal{C}_{B}^{bon}}\right)^{0.5}=0.$ (C.3)
It is straightforward to see that the Left-Hand Side of eq. C.3 is strictly convex and is negative when $Y_{A}=0$. So there is a unique non-negative solution to Equation C.3. Denote this unique solution $Y_{A}^{*}$. Finally, because the term $\mathcal{C}_{A}^{bon}/\mathcal{C}_{B}^{bon}$ only enters negatively, it is also clear that the value of $Y_{A}$ that solves Equation C.3 is strictly increasing in $\mathcal{C}_{A}^{bon}/\mathcal{C}_{B}^{bon}$. ∎
Appendix D Absolute Status
Agents gain status from belonging to a group with high average consumption. An important feature of our model is that group status is relative. What matters is average consumption of an agent’s own group relative to the other group. This notion that status is relative is perhaps the standard view.
Nevertheless, here we show what happens if status depends instead on absolute group consumption. This generates one change in the model: an alternative definition of $Y_{\theta}$. With absolute status, we now have $Y_{\theta}=\mathbf{x}_{\theta}$. This change has very little impact on equilibrium play. All that changes is the definition of $Y^{*}_{\theta_{j}}$. Formally:
Proposition D.1.
$x_{j}^{*}=S^{\prime}_{x}(Y_{\theta_{j}}^{*})\cdot C_{j},\text{ where }Y_{\theta_{j}}^{*}\text{ solves }Y_{\theta}=S^{\prime}_{x}(Y_{\theta})\mathcal{C}_{\theta}$ (D.1)
Proof. Follows directly from the proof to 1
What it does do, however, is shut down the strategic interactions between different groups. Behaviour of agents in group $\theta^{\prime}$ now has no impact on agents in group $\theta$. This induces two changes to the results in Section 3: in 1 $xj^{*}$ is now invariant in $\mathcal{C}_{-\theta}$, and in 2(i) if $\theta_{j}\neq\theta_{k}$ then $x_{j}^{*}$ is invariant in $w_{k}$.
This means that the model with absolute status is unable to explain our third stylised fact (Fact 3). 3(iii) no longer holds with absolute status. Additionally, each group makes decisions independent of all other groups. It is through the relative notion of status that our model is linking the strategic decisions of different groups.